# Error Bounds and Optimal Schedules for Masked Diffusions with Factorized Approximations

### v1

@@ 4.4 The case of a fixed, randomly-generated ordering
Nonetheless, analogous guarantees can be obtained for strategies that first pick a random order (independent of $\pi$), and then keep it fixed during generation. Indeed, since $E_{\text{fact}}=\mathbb{E}_{\nu^{\theta}(\mathbf{z})}[E_{\text{fact}}(\mathbf{z})]$ with $E_{\text{fact}}(\mathbf{z})=\sum_{k\geq 1}\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})\geq 0$, a simple application of Markov’s inequality yields $\Pr_{\nu^{\theta}(\mathbf{z})}(E_{\text{fact}}(\mathbf{z})\geq cE_{\text{fact}})\leq 1/c$, so that choosing $s_{\max}\leq\lceil N/K\rceil$ Theorem 7 implies
@@ 5.3 The case of a diverging number of unmasked variables
$f(i)\approx\frac{1}{N-i}\sum_{j\notin\mathbf{z}}\sum_{\ell\in\mathcal{X}}\log p^{\theta}(x_{j}=\ell;\mathbf{x}_{\mathbf{z}})\,$, (20)

### v2

@@ 4.4 The case of a fixed, randomly-generated ordering
Nonetheless, analogous guarantees can be obtained for strategies that first pick a random order (independent of $\pi$), and then keep it fixed during generation. Indeed, since $E_{\text{fact}}=\mathbb{E}_{\nu^{\theta}(\mathbf{z})}[E_{\text{fact}}(\mathbf{z})]$ with $E_{\text{fact}}(\mathbf{z})=\sum_{k\geq 1}\mathbb{E}_{\pi(\mathbf{x}_{\mathbf{z}_{<k}})}[\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})]\geq 0$, a simple application of Markov’s inequality yields $\Pr_{\nu^{\theta}(\mathbf{z})}(E_{\text{fact}}(\mathbf{z})\geq cE_{\text{fact}})\leq 1/c$, so that choosing $s_{\max}\leq\lceil N/K\rceil$ Theorem 7 implies
@@ 5.3 The case of a diverging number of unmasked variables
$f(i)\approx\frac{1}{N-i}\sum_{j\notin\mathbf{z}}\sum_{\ell\in\mathcal{X}}p^{\theta}(x_{j}=\ell;\mathbf{x}_{\mathbf{z}})\log p^{\theta}(x_{j}=\ell;\mathbf{x}_{\mathbf{z}})\,$, (18)